the botanist decided to count the number of daisies x in each of the 80 randomly selected squares with the field. find the mean and variance


1
Expert's answer
2021-07-01T06:54:35-0400

The results are summarised below


"\sum x=295, \sum x^2=1386"

"Mean=\bar{x}=\dfrac{\sum x}{n}=\dfrac{295}{80}=3.6875\approx3.69"

"Variance=\sigma^2=\dfrac{\sum x^2}{n}-(\bar{x})^2=\dfrac{1386}{80}-(3.6875)^2"

"=\dfrac{1386}{80}-(3.6875)^2=3.72734375\approx3.73"